Grade 10IIT JEE Entrance Exam

This question is from Binomial Theorem Q. Find the value of r, if the coefficients of (2r+4)th and (r-2)th terms in the expansion of (1+x)^18 are equal.

Pranav Pramod Menon
11 Years agoGrade 10

1 Answer

Abhishek Singh
11 Years ago
the binomial cofficients of given terms are 
18C2r+3=18Cr-3
sol.  either 2r+3=r-3          or     2r+3+r-3=18
          no soltion                        r=6